OpenStudy (sleepyjess):

Need help on one question and just want someone to check my answer on another

OpenStudy (sleepyjess):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = \[4/x^2 +9\] f(x) = x + 9 g(x) = \[4/x^2\] f(x) = x g(x)=\[4/x +9\] f(x) = \[1/x\] g(x) = \[4/x +9\] f(x)= \[4/x^2\] g(x) = 9

OpenStudy (sleepyjess):

I think this one is C

OpenStudy (sleepyjess):

f(x) = Square root of quantity three x plus seven. , g(x) = Square root of quantity three x minus seven.. Find (f + g)(x). Square root of six x. <------- x Square root of six. 3x Square root of quantity three x plus seven. +Square root of quantity three minus plus seven.